OpenStudy (anonymous):

Find the slope of the line passing through the points (1)/(9), (-3) (-2), (-1)/(7)

OpenStudy (raden):

can u retype it in order pairs (x1,y1) and (x2,y2) ?

OpenStudy (anonymous):

(1/9,-3) (-2, -1/7)

OpenStudy (raden):

that make more sense :) use the formula : m = (y2-y1)/(x2-x1) with m is the slope so, m = (-1/7 - (-3))/(-2 - 1/9) = ... simplify this

OpenStudy (raden):

did u get it ? @bhow21

OpenStudy (anonymous):

yes thank you could you help me with a few more?

OpenStudy (raden):

sure

OpenStudy (anonymous):

Find the x and y intercepts of the following: (-1)/(8)x-(1)/(3)y=(1)/(12)

OpenStudy (raden):

you mean :